[Brahms-l] Shift report 20050215 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Feb 15 2005 - 08:06:25 EST
Shifter leader was SZ
Shifters were 


1. Beam on  from entry of shift. The RHIC machine worked well!
2. Runs:
13193: tr2=   81k, tr3= 108K  
13194: tr2=  230k, tr3= 125k
13195: tr2=  165k, tr3= 162k
13196: tr2=  115k, tr3= 196k
13197: tr2=  130k, tr3= 124k
13198: tr2=  230k, tr3= 125k
13199: tr2=   65k, tr3= 60k

3. No detector problem

E-log page for this shift:
<a href="http://brahms-web.brahms.bnl.gov/cgi-bin/elog.pl?nb=2005&action=view&page=113&scroll=false">http://brahms-web.brahms.bnl.gov/cgi-bin/elog.pl?nb=2005&action=view&page=113&scroll=false</a>

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 15 08:06:41 2005

This archive was generated by hypermail 2.1.8 : Tue Feb 15 2005 - 08:06:49 EST